OpenStudy (anonymous):

Find the y-intercept for -3x – 2y = 12

OpenStudy (anonymous):

rearrange it as follows -2y = 12 + 3x 12 + 3x y = -------- -2 3 y = - ---- x - 6 2 so putting x = 0, we get y = -6 so the y-intercept is -6